# Modeling extremal dependence in multivariate and spatial problems: a practical perspective

### v3

@@ 2.4.4 Estimation of extreme quantile regions
Theory. As highlighted in Section 2.1, an important problem is to define a quantile region $\mathcal{Q}$ as in (2.11), given a very small probability $p$ to fall in it. Given a sample of size $N$, an extreme quantile region can be defined by the level set $\mathcal{Q}_{N}$ such that ${\mathbb{P}}(\mathcal{Q}_{N})=p$, where $p=p_{N}$ satisfies $p\to 0$ and $Np\to c>0$ as $N\to\infty$. For practical purposes, $\mathcal{Q}_{N}$ can be approximated by the region $\widetilde{\mathcal{Q}}_{N}$ given in (2.13) which requires estimating the marginal parameters ${\bm{\theta}}_{j}=(\mu_{j},\sigma_{j},\gamma_{j})^{\top}$, $j=1,2$, the basic set $S$ and the density of the angular measure, allowing, in turn, to obtain an estimate of $\xi(S)$. Beranger et al. (2021) discussed a Bayesian approach for the inference of $\widetilde{\mathcal{Q}}_{N}$ based on a polynomial angular measure in Bernstein form and the censored-likelihood approach (e.g., Beirlant et al., 2004, Ch. 8). Since such a methodology is similar to the one described in Section “Bernstein polynomials modeling and Bayesian nonparametric inference”, we refer to Beranger et al. (2021) for a full description.
The methodology is illustrated on the dataset Milan.winter, which consists of air pollution levels recorded in Milan, Italy, over the winter period October 31st–February 27/28th, between December 31st 2001 and December 30th 2017. We focus here on the $\textrm{NO}_{2}$ and $\textrm{SO}_{2}$ pollutants and use the daily maximum temperature (MaxTemp) as covariate. The data is prepared before estimating extreme quantile regions using fExtDep.np.
@@
As mentioned above, the estimation is performed through the fExtDep.np function where starting values for the marginal parameters $\bm{\theta}_{j}^{(0)}=\left(\beta_{0,j}^{(0)},\beta_{1,j}^{(0)},\beta_{2,j}^{(0)},\sigma_{j}^{(0)},\gamma_{j}^{(0)}\right)^{\top}$, $j=1,2$, the scaling parameter of the sampler on each margin and the polynomial degree are respectively set by the arguments par10, par20, sig10 sig20 and k0. The argument method = "Bayesian" indicates that a Bernstein polynomial is used to represent the extremal dependence and that the inference part follows the MCMC scheme detailed in Beranger et al. (2021, Algorithm 1). The argument u = TRUE specifies that a censored likelihood approach is applied on raw data with threshold u set to the marginal $90\%$ quantile by default. Recall that by default the marginal distributions are fitted jointly with the dependence (mar.fit = TRUE) but mar.prelim = FALSE indicates that no initial marginal fit is required. The below code chunk provides a full specification of the fExtDep.np arguments.
@@
U set to 90% quantile by default on both marginsEstimation of the extremal dependence and margins |====================================================================| 100%
@@
Figure 11: Top row: posterior mean estimate (dotted line) and 90% credible interval (grey) for the inverse of the angular basic density (left), the basic set $\mathcal{S}$ (middle), and observed data (right) with temperature dependent data colouring (from cold = blue to warm = red). Bottom row: posterior mean estimate (dotted line) and 90% credible interval for the extreme bivariate quantiles associated with probabilities $p=1/600,1/1200$ and $1/2400$ (left to right) for three maximum daily temperature levels: minimum temperature = blue, median temperature = purple and maximum temperature = red.
@@
An implementation is given below where the argument dep=TRUE produces the top row of Figure 11.
@@ 3 Spatial extremes
The fExtDepSpat procedure takes advantage of the availability of the time of occurrence of block maxima to fit the extremal-$t$ (model = "ET") and skew-$t$ (model = "EST") max-stable models using the Stephenson—Tawn likelihood (Stephenson and Tawn, 2005). The correlation function is currently restricted to the power exponential, meaning that the parameters vectors to be estimated are respectively ${\bm{\theta}}=(\nu,r,\eta)^{\top}$ and ${\bm{\theta}}=(\nu,r,\eta,\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$. The corresponding parameters can be fixed through arguments range, smooth, DoF and alpha of fExtDepSpat, else are estimated. Note that for alpha, a vector of length $3$ must be provided and therefore a NA value leaves the corresponding parameter free, e.g., alpha = c(0, NA, NA) would fit the extremal skew-$t$ model with skewness $\bm{\alpha}=\alpha_{1}\textbf{cov}_{1}+\alpha_{2}\textbf{cov}_{2}$. Initial values are provided in par0 in vector form as c(DoF, range, smooth, alpha0, alpha1, alpha2), where fixed or unnecessary parameters are simply omitted. Computations can also be sped-up by considering parallelization (parallel = TRUE) and specifying a number of cores (ncores) to be used. Arguments args1 and args2 are related to specifications of the Monte Carlo simulation scheme to compute the multivariate $t$ cumulative distribution function (cdf). These should take the form of lists including the minimum and maximum number of simulations used (Nmin and Nmax), the absolute error (eps) and whether the error should be controlled on the log-scale (logeps). args1 refers to the d1d-1 dimensional cdfs terms required to compute the exponent function while args2 focuses on the dmd-m ($2\leqslant m\leqslant d-1$) dimensional cdfs involved in the evaluation of its partial derivatives. When computing the log-likelihood function, the latter terms need to be evaluated on the log-scale, requiring fewer Monte Carlo simulations. In the following command lines, the strategy is to set a higher number of simulations for the partial derivative terms since experiments have shown that they are more important than the terms in the exponent function (see Beranger et al., 2021). The control argument offers additional control parameters for the optimization algorithm see ?optim for more details.
@@
$w_{q}=\left\{\begin{array}[]{cl}1&\textrm{if }\max_{i,k\in q;i\neq k}\|s_{i}-s_{k}\|<u\\ 0&\textrm{otherwise}\end{array}.\right$.
@@
Application. We now analyze temperature data around Melbourne, Australia, collected at $90$ stations on a $0.15$ degree (approximately $13$ kilometer) grid in a $9$ by $10$ formation, over the extended summer period from August to April between $1961$ and $2010$. Running the command data(heat, package = "ExtremalDep") loads several important datasets in R. The objects locgrid, scalegrid and shapegrid are matrices of the marginal GEV parameters estimated over the grid using unconstrained location and scale while the shape parameter is defined as a linear function of eastings and northings in $100$ kilometer units. Below, the terra package (Hijmans, 2024) is leveraged to provide graphical illustrations of these objects.
@@
Below the extremal skew-$t$ model with $\nu=5$ is fitted to the data. Since the observations were marginalized with GEV shape parameter as function of eastings and northings in 100 kilometre units, the site locations are transformed to be on the same scale. Note that Beranger et al. (2021) used the TIC for model selection between the extremal-$t$ and extremal skew-$t$ with $\nu=1,3$ and $5$ degrees, but for simplicity we only consider here the selected model and a subset of the locations (ind_sites).
@@ 4 Conclusions
This paper has presented a comprehensive overview of methodological and practical modeling of multivariate and spatial problems. Building on the componentwise maxima framework, we have shown how extremal dependence—characterized through angular measures and Pickands dependence functions—can be effectively estimated, interpreted, and exploited for risk assessment. A key message throughout the paper is that modeling extremal dependence is not only relevant for block maxima, but also provides a powerful basis for approximating probabilities of multiple threshold exceedances, joint tail events, and extreme quantile regions, thereby supporting a broad range of applied risk analyses.
The ExtremalDep package provides a flexible and extensible software that integrates parametric, semiparametric, and nonparametric approaches to modeling extremal dependence. Compared with existing software, it places particular emphasis on nonparametric and semiparametric representations of extremal dependence, to remain close to its nature, while still accommodating parametric models when interpretability or parsimony is required. In addition, it offers a Bayesian inference toolkit that allows for a principled and accessable uncertainty quantification. Through a series of real-world applications—ranging from air pollution and precipitation to financial returns, wind extremes, and heatwaves—the paper has demonstrated how the package can be used to estimate extremal dependence, compute joint and conditional tail probabilities, derive multivariate return levels, simulate extreme events, and construct extreme quantile regions in both multivariate and spatial settings.
Future developments are the following. An important extension concerns the integration of a more a general peaks-over-threshold inferential framework for multivariate data to better complement the componentwise maxima. Scaling nonparametric and Bayesian approaches to even higher dimensions, for example through sparsity assumptions, graphical representations of extremal dependence, or dimension-reduction techniques. In the spatial domain, incorporating spatio-temporal dynamics and more flexible covariate effects into extremal dependence models remains a challenging avenue for research. Future versions of ExtremalDep will aim to improve computational efficiency, for instance by exploiting parallel computing and more efficient Monte Carlo schemes, and to expand the range of implemented models and diagnostic tools and providing additional visualization and decision-oriented outputs.
@@ A.2 Adaptive Gaussian random-walk Metropolis–Hastings (RWMH) algorithm
where $c=(2\pi)^{1/2}\exp\left(\zeta_{0}^{2}/2\right)/(2\zeta_{0})$ is a steplength constant, $\zeta_{0}=-1/\Phi(\pi^{*}/2)$, and $\Phi$ is the univariate standard Gaussian distribution function (Garthwaite et al., 2016). To control the desired sampler acceptance probability, the parameter $\pi^{*}=0.234$ is specified as in Roberts et al. (1997).
@@ 3      Step 1: Marginal component 1:
4       Draw proposal ${\bm{\theta}}_{1}^{\prime}\sim MVN({\bm{\theta}}_{1}^{(s)},\tau^{(s)}_{1}\Sigma^{(s)}_{1})$.
5       Compute acceptance probability $\pi_{1}=\min\left\{\frac{{\mathcal{L}}\left({\bm{\theta}}_{1}^{\prime},{\bm{\theta}}_{2}^{(s)},\kappa^{(s)},{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi({\bm{\theta}}_{1}^{\prime})}{{\mathcal{L}}\left({\bm{\theta}}_{1}^{(s)},{\bm{\theta}}_{2}^{(s)},\kappa^{(s)},{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi({\bm{\theta}}_{1}^{(s)})},1\right\}.$
@@ 9      Step 2: Marginal component 2:
10       Draw proposal ${\bm{\theta}}_{2}^{\prime}\sim MVN({\bm{\theta}}_{2}^{(s)},\tau^{(s)}_{2}\Sigma^{(s)}_{2})$.
11       Compute acceptance probability $\pi_{2}=\min\left\{\frac{{\mathcal{L}}\left({\bm{\theta}}_{1}^{(s+1)},{\bm{\theta}}_{2}^{\prime},\kappa^{(s)},{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi({\bm{\theta}}_{2}^{\prime})}{{\mathcal{L}}\left({\bm{\theta}}_{1}^{(s+1)},{\bm{\theta}}_{2}^{(s)},\kappa^{(s)},{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi({\bm{\theta}}_{2}^{(s)})},1\right\}.$

### v5

@@ 2.4.4 Estimation of extreme quantile regions
Theory. As highlighted in Section 2.1, an important problem is to define a quantile region $\mathcal{Q}$ as in (2.11), given a very small probability $p$ to fall in it. Given a sample of size $N$, an extreme quantile region can be defined by the level set $\mathcal{Q}_{N}$ such that ${\mathbb{P}}(\mathcal{Q}_{N})=p$, where $p=p_{N}$ satisfies $p\to 0$ and $Np\to c>0$ as $N\to\infty$. For practical purposes, $\mathcal{Q}_{N}$ can be approximated by the region $\widetilde{\mathcal{Q}}_{N}$ given in (2.13) which requires estimating the marginal parameters ${\bm{\theta}}_{j}=(\mu_{j},\sigma_{j},\gamma_{j})^{\top}$, $j=1,2$, the basic set $S$ and the density of the angular measure, allowing, in turn, to obtain an estimate of $\xi(S)$. Beranger et al. (2021) discussed a Bayesian approach for the inference of $\widetilde{\mathcal{Q}}_{N}$ based on a polynomial angular measure in Bernstein form and the censored-likelihood approach (e.g., Beirlant et al., 2004, Ch. 8). Since such a methodology is similar to the one described in Section 2.4.2, we refer to Beranger et al. (2021) for a full description.
The methodology is illustrated on the dataset Milan.winter, which consists of air pollution levels recorded in Milan, Italy, over the winter period October 31st–February 28/29th, between December 31st 2001 and December 30th 2017. We focus here on the $\textrm{NO}_{2}$ and $\textrm{SO}_{2}$ pollutants and use the daily maximum temperature (MaxTemp) as covariate. The data is prepared before estimating extreme quantile regions using fExtDep.np.
@@
As mentioned above, the estimation is performed through the fExtDep.np function where starting values for the marginal parameters $\bm{\theta}_{j}^{(0)}=\left(\beta_{0,j}^{(0)},\beta_{1,j}^{(0)},\beta_{2,j}^{(0)},\sigma_{j}^{(0)},\gamma_{j}^{(0)}\right)^{\top}$, $j=1,2$, the scaling parameter of the sampler on each margin and the polynomial degree are respectively set by the arguments par10, par20, sig10 sig20 and k0. The argument method = "Bayesian" indicates that a Bernstein polynomial is used to represent the extremal dependence and that the inference part follows the MCMC scheme detailed in Beranger et al. (2021, Algorithm 1). The argument u = TRUE specifies that a censored likelihood approach is applied on raw data with threshold u set to the marginal $90\%$ quantile by default. Recall that by default the marginal distributions are fitted jointly with the dependence (mar.fit = TRUE) but mar.prelim = FALSE indicates that no initial marginal fit is required. The code chunk below provides a full specification of the fExtDep.np arguments.
@@
Figure 11: Top row: posterior mean estimate (dotted line) and 90% credible interval (grey) for the inverse of the angular basic density (left), the basic set $\mathcal{S}$ (middle), and observed data (right) with temperature dependent data colouring (from cold = blue to warm = red). Bottom row: posterior mean estimate (dotted line) and 90% credible interval for the extreme bivariate quantiles associated with probabilities $p=1/600,1/1200$ and $1/2400$ (left to right) for three maximum daily temperature levels: minimum temperature (blue), median (purple) and maximum (red) temperatures.
@@
An implementation is given below, where the argument dep=TRUE produces the top row of Figure 11.
@@ 3 Spatial extremes
The fExtDepSpat procedure takes advantage of the availability of the time of occurrence of block maxima to fit the extremal-$t$ (model = "ET") and skew-$t$ (model = "EST") max-stable models using the Stephenson—Tawn likelihood (Stephenson and Tawn, 2005). The correlation function is currently restricted to the power exponential, meaning that the parameters vectors to be estimated are respectively ${\bm{\theta}}=(\nu,r,\eta)^{\top}$ and ${\bm{\theta}}=(\nu,r,\eta,\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$. The corresponding parameters can be fixed through arguments range, smooth, DoF and alpha of fExtDepSpat, else are estimated. Note that, a vector of length $3$ must be provided for alpha and therefore a NA value leaves the corresponding parameter free, e.g., alpha = c(0, NA, NA) would fit the extremal skew-$t$ model with skewness $\bm{\alpha}=\alpha_{1}\textbf{cov}_{1}+\alpha_{2}\textbf{cov}_{2}$. Initial values are provided in par0 in vector form as c(DoF, range, smooth, alpha0, alpha1, alpha2), where fixed or unnecessary parameters are simply omitted. Computations can also be sped-up by considering parallelization (parallel = TRUE) and specifying the number of cores (ncores) to be used. Arguments args1 and args2 are related to specifications of the Monte Carlo simulation scheme to compute the multivariate $t$ cumulative distribution function (cdf). These should take the form of lists including the minimum and maximum number of simulations used (Nmin and Nmax), the absolute error (eps) and whether the error should be controlled on the log-scale (logeps). args1 refers to the d1d-1 dimensional cdfs terms required to compute the exponent function while args2 focuses on the dmd-m ($2\leqslant m\leqslant d-1$) dimensional cdfs involved in the evaluation of its partial derivatives. When computing the log-likelihood function, the latter terms need to be evaluated on the log-scale, requiring fewer Monte Carlo simulations. In the following command lines, the strategy is to set a higher number of simulations for the partial derivative terms since experiments have shown that they are more important than the terms in the exponent function (see Beranger et al., 2021). The control argument offers additional control parameters for the optimization algorithm see ?optim for more details.
@@
$w_{q}=\left\{\begin{array}[]{cl}1&\textrm{if }\max_{i,k\in q;i\neq k}\|s_{i}-s_{k}\|<u,\\ 0&\textrm{otherwise}.\end{array}\right$.
@@
Application. We now analyze temperature data around Melbourne, Australia, collected at $90$ stations on a $0.15$ degree (approximately $13$ kilometer) grid in a $9$ by $10$ formation, over the extended summer period from August to April between $1961$ and $2010$. Running the command data(heat, package = "ExtremalDep") loads several datasets in R. The objects locgrid, scalegrid and shapegrid are matrices of the marginal GEV parameters estimated over the grid using unconstrained location and scale while the shape parameter is defined as a linear function of eastings and northings in $100$ kilometer units. Below, the terra package (Hijmans, 2026) is leveraged to provide graphical illustrations of these objects.
@@
Below, we fit the extremal skew-$t$ model with $\nu=5$ to the data. Since the observations were marginalized with GEV shape parameter as function of eastings and northings in 100 kilometre units, the site locations are transformed to be on the same scale. Beranger et al. (2021) used the TIC for model selection between the extremal-$t$ and extremal skew-$t$ with $\nu=1,3$ and $5$ degrees but, for simplicity, we only consider here the selected model and a subset of the locations (ind_sites).
@@ 4 Conclusions
This paper has presented a comprehensive overview of methodological and practical modeling of multivariate and spatial problems. Building on the componentwise maxima framework, we have shown how extremal dependence, characterized through angular measures and Pickands dependence functions, can be effectively estimated, interpreted, and exploited for risk assessment. A key message throughout the paper is that modeling extremal dependence is not only relevant for block maxima, but also provides a powerful basis for approximating probabilities of multiple threshold exceedances, joint tail events, and extreme quantile regions, thereby supporting a broad range of applied risk analyses.
The ExtremalDep package provides a flexible and extensible software that integrates parametric, semiparametric, and nonparametric approaches to modeling extremal dependence. Compared with existing software, it places particular emphasis on nonparametric and semiparametric representations of extremal dependence, to remain close to its nature, while still accommodating parametric models when interpretability or parsimony is required. In addition, it offers a Bayesian inference toolkit that allows for a principled and accessable uncertainty quantification. Through a series of real-world applications, ranging from air pollution and precipitation to financial returns, wind extremes, and heatwaves, the paper has demonstrated how the R package can be used to estimate extremal dependence, compute joint and conditional tail probabilities, derive multivariate return levels, simulate extreme events, and construct extreme quantile regions in both multivariate and spatial settings.
We plan the following future developments. An important extension concerns the integration of a more a general peaks-over-threshold inferential framework for multivariate data to better complement the componentwise maxima. Scaling nonparametric and Bayesian approaches to even higher dimensions, for example through sparsity assumptions, graphical representations of extremal dependence, or dimension-reduction techniques. In the spatial domain, incorporating spatio-temporal dynamics and more flexible covariate effects into extremal dependence models remains a challenging avenue for research. Future versions of ExtremalDep will aim to improve computational efficiency, for instance by exploiting parallel computing and more efficient Monte Carlo schemes, and to expand the range of implemented models and diagnostic tools and providing additional visualization and decision-oriented outputs.
@@ A.2 Adaptive Gaussian random-walk Metropolis–Hastings (RWMH) algorithm
where $c=(2\pi)^{1/2}\exp\left(\zeta_{0}^{2}/2\right)/(2\zeta_{0})$ is a steplength constant, $\zeta_{0}=-1/\Phi(\pi^{*}/2)$, and $\Phi$ is the univariate standard Gaussian distribution function (Garthwaite et al., 2016). To control the desired sampler acceptance probability, the parameter $\pi^{*}=0.234$ is specified as in Roberts et al. (1997). We note that some improvements of the adaptive MCMC could be implemented to gain in efficiency, for example by taking as initial covariance a multiple of the inverse Fisher information.
@@ 3      Step 1: Marginal component 1:
4       Draw proposal ${\bm{\theta}}_{1}^{\prime}\sim\mathrm{MVN}({\bm{\theta}}_{1}^{(s)},\tau^{(s)}_{1}\Sigma^{(s)}_{1})$.
5       Compute acceptance probability $\pi_{1}=\min\left\{\frac{{\mathcal{L}}\left({\bm{\theta}}_{1}^{\prime},\,{\bm{\theta}}_{2}^{(s)},\,\kappa^{(s)},\,{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi({\bm{\theta}}_{1}^{\prime})}{{\mathcal{L}}\left({\bm{\theta}}_{1}^{(s)},\,{\bm{\theta}}_{2}^{(s)},\,\kappa^{(s)},\,{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi\left({\bm{\theta}}_{1}^{(s)}\right)},1\right\}.$
@@ 9      Step 2: Marginal component 2:
10       Draw proposal ${\bm{\theta}}_{2}^{\prime}\sim\mathrm{MVN}({\bm{\theta}}_{2}^{(s)},\tau^{(s)}_{2}\Sigma^{(s)}_{2})$.
11       Compute acceptance probability $\pi_{2}=\min\left\{\frac{{\mathcal{L}}\left({\bm{\theta}}_{1}^{(s+1)},\,{\bm{\theta}}_{2}^{\prime},\,\kappa^{(s)},\,{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi({\bm{\theta}}_{2}^{\prime})}{{\mathcal{L}}\left({\bm{\theta}}_{1}^{(s+1)},\,{\bm{\theta}}_{2}^{(s)},\,\kappa^{(s)},\,{\bm{\beta}}_{\kappa^{(s)}}^{(s)}\right)\Pi\left({\bm{\theta}}_{2}^{(s)}\right)},1\right\}.$